(1 point) The gender and enrollment status of one-hundred students is summarized in the table below.

Female Male Total
Part-time 2323 2828 51
Full-time 1313 3636 4949
Total 3636 6464 100100


If a student is selected at random from those surveyed, find the probability of each of the following events.

1. The student is female or is enrolled part-time.
Answer:

2. The student is male or is enrolled full-time.
Answer:

3. The student is female or is enrolled full-time.
Answer:
Answer #1

Solution :

From given table,

1)

P(female or Part-time)

P(female) + P(Part-time) - P(female or Part-time)

= (36 + 51 - 23) / 100

= 64 / 100

= 0.64

2)

P(male or Part-time)

= (64 + 51 - 28) / 100

= 87 / 100

= .87

3)

P(female or Full-time)

= (36 + 49 - 13) / 100

= 72 / 100

= 0.72